# RealGeneral: Unifying Visual Generation via Temporal In-Context Learning with Video Models

### v1

@@ Abstract
*Corresponding author
Unifying diverse image generation tasks within a single framework remains a fundamental challenge in visual generation. While large language models (LLMs) achieve unification through task-agnostic data and generation, existing visual generation models fail to meet these principles. Current approaches either rely on per-task datasets and large-scale training or adapt pre-trained image models with task-specific modifications, limiting their generalizability. In this work, we explore video models as a foundation for unified image generation, leveraging their inherent ability to model temporal correlations. We introduce RealGeneral, a novel framework that reformulates image generation as a conditional frame prediction task, analogous to in-context learning in LLMs. To bridge the gap between video models and condition-image pairs, we propose (1) a Unified Conditional Embedding module for multi-modal alignment and (2) a Unified Stream DiT Block with decoupled adaptive LayerNorm and attention mask to mitigate cross-modal interference. RealGeneral demonstrates effectiveness in multiple important visual generation tasks, e.g., it achieves a 14.5% improvement in subject similarity for customized generation and a 10% enhancement in image quality for canny-to-image task. Project page: https://lyne1.github.io/RealGeneral/
Figure 1: Results from our RealGeneral, demonstrate the ability to produce high-quality images from diverse input conditions.
@@ 1 Introduction
With the rapid development of generative modeling [15, 16, 5, 20, 51, 26], image generation has gained considerable attention in various application domains. Although foundation models such as Stable Diffusion [10, 32] and FLUX [22] have achieved remarkable success in unimodal text-to-image generation, processing diverse inputs (e.g., text, depth maps, conditional images) and generating corresponding images within a unified framework remains a significant challenge. While large language models (LLMs) have unified text generation [4, 48, 1], an analogous unified model for the visual domain is still lacking.
Upon closer examination of the core principles underlying LLMs’ success as a unified text generation architecture, we identify two fundamental pillars, i.e., (1) a task-agnostic data principle during pre-training that facilitates the utilization of web-scale corpora for general linguistic knowledge acquisition, and (2) a task-agnostic generation principle that coherently unifies all kinds of textual tasks through next-token prediction. Analogously, a promising and effective architecture for unified visual generation should also exhibit (1) a task-agnostic visual data principle for general visual knowledge acquisition, and (2) a task-agnostic visual generation principle that could unify all kinds of multi-modal visual tasks (e.g., customized generation, canny-to-image generation, etc.), as shown in Fig. 2.
@@ 2.1 Image Generation
Image generation is an area of rapid development that integrates concepts from natural language processing with advancements in transformer architectures. Autoregressive models utilize transformers to predict sequences of discrete codebook codes [10, 35, 40]. Meanwhile, continuous diffusion models have emerged as a powerful framework for image generation [15, 37, 9]. The integration of large-scale transformer architectures has led to the development of advanced models, such as DiT [31, 22]. Many efforts have been made to extend the capabilities of generation models, with notable techniques, such as T2I-Adapter [28]. However, these methods are designed for specific tasks, enhancing the capabilities of generative models through architectural modifications.
@@ 2.3 Unified Image Generation
Current research on unified image generation can be mainly categorized into two streams: training-from-scratch, and image-model-adaptation. Many works focus on the first stream [33, 45, 27, 2, 42, 8, 13], For example, PixWizard [25] employs a flow-based DiT [31] model that gradually injects conditional and textual information during cross-attention while computing only the most relevant semantic tokens. OmniGen [50] jointly models text and images, applying causal attention to each token in the sequence and bidirectional attention within each image sequence. The image-model-adaptation stream leverages pre-trained image models to improve efficiency. For instance, ControlNet [55] uses trainable copies and zero convolution layers, enabling Stable Diffusion [10, 32] to adapt specific tasks. OminiControl [43] builts on FLUX.1 [22], proposing a multi-modal attention mechanism that facilitates direct interaction between conditions and images.
Our concurrent work, UniReal [8], similarly reformulates images as discrete frames within a video generation framework as we do. However, there are three critical distinctions between RealGeneral and UniReal. First, the motivations diverge: UniReal adopts the training-from-scratch paradigm only modeling the relationship of discrete frames, whereas we propose two visual principles underlying the unification of visual generation, recasting multi-condition image generation as temporal in-context learning within video models, analogous to LLMs. Second, the design strategies are distinct. UniReal relies on complex prompt engineering, which involves multiple special tokens, contextual prompts, and hierarchical prompts, while our method doesn’t need these. Third, our method is markedly more efficient: UniReal necessitates 360 million training samples and substantial computational resources. However, our method fine-tunes a pre-trained video model using LoRA with only 0.1% of that data, significantly reducing the computational cost.
@@ 3.2 Unified Conditional Embedding Module
In various visual generative tasks, inputs are often multi-modal(e.g., text and image), while outputs are image modality. Therefore, the key to unified visual generation lies in how to effectively fuse multi-modal conditions, and avoid confusing conditional and generated images with each other. In RealGeneral, we design the UCE module to align multi-modal conditions, containing two complementary pathways, i.e., the Subject-Specific Embedding layer for inter-modal conditional semantic alignment and the Condition Embedding layer for intra-modal generative semantic distinction.
@@
Before patchify operation, the vision sequence $\mathbf{z}_{vision}\in\mathbb{R}^{2\times h\times w\times d}$ is formed by concatenating $\mathbf{c}_{cond}$ and $\mathbf{x}_{target}$. Patchify operation in CogVideoX divides the input sequence along spatial and temporal dimensions using patch factors $p=2$. This process compresses the temporal dimension, leading to the intermingling of the latent tokens for $\mathbf{z}_{cond}$ and $\mathbf{z}_{target}$. Such mixing obscures the distinct temporal features inherent to each frame. To mitigate this, we propose a simple replication strategy, i.e., we replicate the latent tokens for both $\mathbf{c}_{cond}$ and $\mathbf{x}_{target}$ along the temporal dimension. Finally, $\mathbf{c}_{cond}$, $\mathbf{x}_{target}$ are concatenated into the input sequence for the Unified Stream DiT Block:
@@ 3.3 Separated Condition AdaLN Module
We introduce SC-AdaLN module, a novel mechanism to address the inherent conflict between multi-modal conditions in image generation tasks. Existing video generation frameworks employ two AdaLN modules to handle text and video inputs separately. This leads to interference between the condition frame and target frame due to their shared video AdaLN parameters. Our SC-AdaLN module resolves this by decoupling this process into three independent branches, explicitly disentangling three distinct modalities:
@@ 3.4 Frame-Condition Decoupling Module
We propose the FCD module to mitigate multi-modal condition confusion, mainly containing the Frame-Condition Decoupling Attention (FCD Attention) which addresses the cross-modal condition interference.
@@ 4.2 Main Results
Table 1: Quantitative results with existing method on customized image generation task. We highlight the best and second-best values for each metric.
@@
The quantitative results are presented in Tab. 2. For the canny-to-image task, RealGeneral attains superior FID and SSIM scores compared to existing models. For the depth-to-image task, RealGeneral achieves a balanced performance in terms of FID and SSIM, attaining the second-best for both metrics. In addition, because the depth and canny map contain spatial information, the CLIP-T of the generated images are very close (both 0.328). We don’t show this metric in the table.
@@ 4.3 Base Model Performance
Figure 7: Comparison of image generation results between CogVideoX1.5 and other image base models.
@@ 5 Conclusion & Future Work
In this paper, we propose RealGeneral, a novel framework that bridges the gap between video generation models and unified visual synthesis tasks. we unlock the inherent potential of pre-trained video models for diverse image generation tasks, by reformulating conditional image generation as temporal in-context learning through temporal modeling. This is achieved through three core modules: (1) UCE module that aligns multi-modal semantics, (2) SC-AdaLN that disentangles feature modulation across text, conditional frames, and target frames, (3) FCD module that prevents conditional semantic interference. Together, these components form a solution for unified image generation tasks.

### v2

@@ Abstract
Zhendong Mao is the corresponding author.
Unifying diverse image generation tasks within a single framework remains a fundamental challenge in visual generation. While large language models (LLMs) achieve unification through task-agnostic data and generation, existing visual generation models fail to meet these principles. Current approaches either rely on per-task datasets and large-scale training or adapt pre-trained image models with task-specific modifications, limiting their generalizability. In this work, we explore video models as a foundation for unified image generation, leveraging their inherent ability to model temporal correlations. We introduce RealGeneral, a novel framework that reformulates image generation as a conditional frame prediction task, analogous to in-context learning in LLMs. To bridge the gap between video models and condition-image pairs, we propose (1) a Unified Conditional Embedding module for multi-modal alignment and (2) a Unified Stream DiT Block with decoupled adaptive LayerNorm and attention mask to mitigate cross-modal interference. RealGeneral demonstrates effectiveness in multiple important visual generation tasks, e.g., it achieves a 14.5% improvement in subject similarity for customized generation and a 10% enhancement in image quality for canny-to-image task. Project Page: realgeneral_web; GitHub Link: https://github.com/Lyne1/RealGeneral
Figure 1: Results from RealGeneral, demonstrate the ability to produce high-quality images from diverse input conditions.
@@ 1 Introduction
With the rapid development of generative modeling [15, 16, 20, 5, 21, 54, 28], image generation has gained considerable attention in various application domains. Although foundation models such as Stable Diffusion [10, 35] and FLUX [24] have achieved remarkable success in unimodal text-to-image generation, processing diverse inputs (e.g., text, depth maps, conditional images) and generating corresponding images within a unified framework remains a significant challenge. While large language models (LLMs) have unified text generation [4, 51, 1], an analogous unified model for the visual domain is still lacking.
Upon closer examination of the core principles underlying LLMs’ success as a unified text generation architecture, we identify two fundamental pillars, i.e., (1) a task-agnostic data principle during pre-training that facilitates the utilization of web-scale corpora for general linguistic knowledge acquisition, and (2) a task-agnostic generation principle that coherently unifies all kinds of textual tasks through next-token prediction. Analogously, a promising and effective architecture for unified visual generation should also exhibit (1) a task-agnostic visual data principle for general visual knowledge acquisition, and (2) a task-agnostic visual generation principle that could unify all kinds of multi-modal visual tasks (e.g., customization generation, canny-to-image generation, etc.), as shown in Fig. 2.
@@ 2.1 Image Generation
Image generation has rapidly advanced by combining NLP techniques with transformer architectures. Autoregressive models utilize transformers to predict sequences of discrete codebook codes [10, 38, 43]. Meanwhile, continuous diffusion models have emerged as a powerful framework for image generation [15, 40, 9]. The integration of large-scale transformer architectures has led to the development of advanced models, such as DiT [34, 24]. Many efforts have been made to extend the capabilities of generation models. For example, T2I-Adapter [31] introduces additional encoding layers to improve controllability, while RealCustom [22, 29] disentangles subject similarity from textual controllability by restricting subject influence to semantically relevant regions through a train-inference decoupled framework. However, these methods are designed for specific tasks, enhancing the capabilities of generative models through architectural modifications.
@@ 2.3 Unified Image Generation
Current research on unified image generation can be mainly categorized into two streams: training-from-scratch, and image-model-adaptation. Many works focus on the first stream [36, 48, 30, 2, 45, 8, 13]. For example, PixWizard [27] employs a flow-based DiT [34] model that gradually injects conditional and textual information during cross-attention. OmniGen [53] jointly models text and images, applying causal attention to each token in the sequence and bidirectional attention within each image sequence. The image-model-adaptation stream leverages pre-trained image models to improve efficiency. For instance, ControlNet [58] uses trainable copies and zero convolution layers, enabling Stable Diffusion [10, 35] to adapt specific tasks. OminiControl [46] builds on FLUX.1 [24], proposing a multi-modal attention mechanism that facilitates direct interaction between conditions and images.
Our concurrent work, UniReal [8], also reformulates images as discrete frames within a video generation framework. However, there are three critical distinctions between UniReal and our approach. First, the motivations differ: UniReal adopts the training-from-scratch paradigm, only modeling the relationship between discrete frames, whereas RealGeneral introduces two visual principles underlying the unification of visual generation, recasting multi-condition image generation as temporal in-context learning within video models. Second, the design choices diverge: UniReal relies on complex prompt engineering, while RealGeneral achieves controllability without such designs. Third, RealGeneral is significantly more efficient, fine-tuning a pre-trained video model using LoRA with only 0.1% of UniReal’s 360 million training samples, resulting in substantially reduced computational cost.
@@ 3.2 Unified Conditional Embedding Module
In various visual generative tasks, inputs are often multi-modal(e.g., text and image), while output is image modality. Therefore, the key to unified visual generation lies in how to effectively fuse multi-modal conditions and avoid confusing conditional and generated images with each other. In RealGeneral, we design the UCE module to align multi-modal conditions, containing two complementary pathways, i.e., the Subject-Specific Embedding layer for inter-modal conditional semantic alignment and the Condition Embedding layer for intra-modal generative semantic distinction.
@@
Before patchify operation, the vision sequence $\mathbf{z}_{vision}\in\mathbb{R}^{2\times h\times w\times d}$ is formed by concatenating $\mathbf{c}_{cond}$ and $\mathbf{x}_{target}$. The patchify operation in CogVideoX divides the input sequence along spatial and temporal dimensions using patch factors $p=2$. This process compresses the temporal dimension, leading to the intermingling of the latent tokens for $\mathbf{z}_{cond}$ and $\mathbf{z}_{target}$. Such mixing obscures the distinct temporal features inherent to each frame. To mitigate this, we propose a simple replication strategy, i.e., we replicate the latent tokens for both $\mathbf{c}_{cond}$ and $\mathbf{x}_{target}$ along the temporal dimension. Finally, $\mathbf{c}_{cond}$, $\mathbf{x}_{target}$ are concatenated into the input sequence for the Unified Stream DiT Block:
@@ 3.3 Separated Condition AdaLN Module
We introduce the SC-AdaLN module, a novel mechanism to address the inherent conflict between multi-modal conditions in image generation tasks. Existing video generation frameworks employ two AdaLN modules to handle text and video inputs separately. This leads to interference between the condition frame and target frame due to their shared video AdaLN parameters. Our SC-AdaLN module resolves this by decoupling this process into three independent branches, explicitly disentangling three distinct modalities:
@@ 3.4 Frame-Condition Decoupling Module
We propose the FCD module to mitigate multi-modal condition confusion, mainly containing the Frame-Condition Decoupling Attention (FCD Attention), which addresses the cross-modal condition interference.
@@ 4.2 Main Results
Table 1: Quantitative results with existing methods on customization task. We highlight the best and second-best values for each metric.
@@
The quantitative results are presented in Tab. 2. For the canny-to-image task, RealGeneral attains superior FID and SSIM scores compared to existing models. For the depth-to-image task, RealGeneral achieves a balanced performance in terms of FID and SSIM, attaining the second-best for both metrics. In addition, because the depth and canny maps contain spatial information, the CLIP-T of the generated images is very close (both 0.328). We don’t show this metric in the table.
@@ 4.3 Base Model Performance
Figure 7: Comparison of image generation results between CogVideoX1.5 and other image generation models.
@@ 5 Conclusion & Future Work
In this paper, we propose RealGeneral, a novel framework that bridges the gap between video generation models and unified visual synthesis tasks. We unlock the inherent potential of pre-trained video models for diverse image generation tasks by reformulating conditional image generation as temporal in-context learning through temporal modeling. This is achieved through three core modules: (1) UCE module that aligns multi-modal semantics, (2) SC-AdaLN that disentangles feature modulation across text, conditional frames, and target frames, (3) FCD module that prevents conditional semantic interference. Together, these components form a solution for unified image generation tasks.
@@
Acknowledgment
This research is supported by Artificial Intelligence National Science and Technology Major Project 2023ZD0121200, and National Natural Science Foundation of China under Grant 62222212 and 623B2094.